OpenStudy (anonymous):

Solve simutaneously : (5/2)+(x/8)=y 11+(y/3)=3x

OpenStudy (anonymous):

you have what y equals 3 [ 11 + (5/2+x/8) / 3 = 3x ] get rid of denominator 1/3 multiply by reciprocal 33 + 5/2 + x / 8 = 9 x 33 + 5/2 + x/ 8 +(- x/8 ) = 9 x + ( - x/ 8) [ 33+ 5/2 = 9x - x / 8 ] 2 66 + 5 = 18x - x / 4 264+ 20 = 72x - x 284= 71x 4 = x

OpenStudy (anonymous):

ok i think